# Haojun Chen    Minghao Liu    Xiaojian Ma    Zailin Ma    Huimin Wu    Chengdong Ma    Yuanpei Chen    Yifan Zhong    Mingzhi Wang    Qing Li    Yaodong Yang

### v1

@@ Abstract
Diffusion policies are widely adopted in complex visuomotor tasks for their ability to capture multimodal action distributions. However, the multiple sampling steps required for action generation significantly harm real-time inference efficiency which limits their applicability in long-horizon tasks and real-time decision-making scenarios. Existing acceleration techniques reduce sampling steps by approximating the original denoising process but inevitably introduce unacceptable performance loss. Here we propose Falcon, which mitigates this trade-off and achieves further acceleration. The core insight is that visuomotor tasks exhibit sequential dependencies between actions at consecutive time steps. Falcon leverages this property to avoid denoising from a standard normal distribution at each decision step. Instead, it starts denoising from partial denoised actions derived from historical information to significantly reduce the denoising steps, while incorporating current observations to achieve performance-preserving acceleration of action generation. Importantly, Falcon is a training-free algorithm that can be applied as a plug-in to further improve decision efficiency on top of existing acceleration techniques. We validated Falcon in 46 simulated environments, demonstrating a 2-7x speedup with negligible performance degradation, offering a promising direction for efficient visuomotor policy design.
@@ 1 Introduction
Diffusion policies have demonstrated remarkable success in addressing complex visuomotor tasks in robotics (Chi et al., 2023; Reuss et al., 2023; Ze et al., 2024b, a; Ravan et al., 2024; Yang et al., 2024), thanks to their ability to model complex multimodal distributions and maintain stable training dynamics. Essentially, diffusion policies rely on the reverse sampling process of a stochastic differential equation (SDE) (Song et al., 2020b; Karras et al., 2022), where actions are iteratively sampled starting from a standard normal distribution. Each sampling step involves drawing a sample from a Brownian motion distribution, incrementally denoising the initial sample to generate the final action. However, the iterative sampling process required for action generation makes these methods computationally slow, particularly when applied to sequential decision-making in real-time environments (Chen et al., 2023b; Janner et al., 2022; Wang et al., 2022; Yang et al., 2023; Hansen-Estruch et al., 2023; Chen et al., 2023a).
To address these challenges, existing work (Song et al., 2020a, b; Lu et al., 2022; Zhang & Chen, 2022) reformulates the sampling process as an ordinary differential equation (ODE), which allows the use of numerical solvers to approximate the ODE’s solution and reduce the number of required sampling steps. However, using numerical solvers introduces inherent approximation errors, which can prevent the generation of high-quality actions in a few steps (Zhao et al., 2024). Additionally, some methods (Song et al., 2023; Kim et al., 2023; Prasad et al., 2024; Wang et al., 2024) employ distillation techniques to achieve single-step generation. While distillation improves efficiency, it compromises the ability of diffusion policies to represent multimodal action distributions, resulting in significant performance losses. Moreover, distillation-based approaches are inherently training-intensive and task-specific, meaning they cannot generalize effectively to accelerate unseen tasks or adapt to diverse visuomotor applications.
@@
In this work, we introduce Falcon (Fast visuomotor policies via partial denoising), a novel approach designed to bridge the gap between acceleration and performance preservation in diffusion-based visuomotor policies. The key insight behind Falcon is that sequential dependencies in visuomotor tasks can be exploited to accelerate action generation while maintaining multimodal expressiveness. To effectively leverage this property, we first utilize the previous action prediction as the reference action, then introduce a thresholding mechanism combined with one-step estimation to evaluate which partial denoising actions are dependent on current timesteps in parallel. Using a temperature-scaled softmax, we then select the most suitable partial denoising action to continue the sampling process, preserving both performance and efficiency. By avoiding the conventional practice of starting the denoising process from a standard normal distribution at each decision step, Falcon begins from partial denoised actions derived from historical observations, significantly reducing the number of sampling steps required. Importantly, Falcon is a training-free method, which allows it to be applied as a plug-in module, enhancing the efficiency of existing diffusion-based policies without additional training or extensive modifications. It integrates seamlessly with solvers like DDIM (Song et al., 2020a) and DPMSolver (Lu et al., 2022) to further enhance acceleration.
We evaluate Falcon on various simulated environments, including RoboMimic (Mandlekar et al., 2021), RoboSuite Kitchen (Gupta et al., 2019), BlockPush (Shafiullah et al., 2022) and MetaWorld (Yu et al., 2020). The results demonstrate that Falcon achieves a 2-7× speedup in action generation with negligible performance degradation. By mitigating the trade-off between speed and performance, Falcon offers a promising direction for advancing the capabilities of diffusion policies in both simulated and real-world settings.
@@
In summary, this work makes three main contributions: First, we introduce Falcon, which leverages sequential dependencies in visuomotor tasks to perform partial denoising, significantly reducing the number of sampling steps while preserving the ability to model multimodal action distributions. Second, Falcon functions as a training-free plug-in that enhances existing diffusion-based policies, integrating seamlessly with solvers like DDIM and DPMSolver to further accelerate action generation. Third, we demonstrate Falcon’s effectiveness across 46 simulated environments, achieving a 2-7× speedup in inference while maintaining high action quality.
@@ 2 Preliminaries
In this section, we give a brief introduction to the diffusion model, the acceleration techniques in the diffusion model, how the diffusion models are used for diffusion policies in visuomotor tasks, and some necessary terminology.
Diffusion Models.   Diffusion models, such as Denoising Diffusion Probabilistic Models (DDPM)(Ho et al., 2020), are generative models that learn data distributions by progressively corrupting data through noise in a forward process and then reconstructing it in a reverse denoising process. In the forward process, a data point $\boldsymbol{x}_{0}$ is corrupted K timesteps, resulting in:
@@
where $\mu_{\theta}$ is the predicted mean, and $\sigma_{t}^{2}$ is fixed according to the forward process.
Accelerate techniques in the diffusion model.   Some popular works (Song et al., 2020b; Karras et al., 2022; Song et al., 2020a; Lu et al., 2022) interpret the diffusion model as an ODE. Specifically, DDIM (Song et al., 2020a) generalizes the Markov forward process in DDPM to a non-Markov process, allowing the use of shorter Markov chains during sampling. The iterative steps of its sampling process can be rewritten in a form similar to Euler integration, which is a discrete solution process for a specific ODE. DPMSolver (Lu et al., 2022) accelerates sampling by analytically computing the linear part of the diffusion ODE solution and using an exponential integrator to approximate the nonlinear part. But these methods can reduce the quality degradation of few-step sampling (Shih et al., 2024).
Diffusion Policies.   Diffusion policy (Chi et al., 2023) extends diffusion models as a powerful policy for visuomotor tasks. At timestep $t$, diffusion policy takes the latest $T_{o}$ steps of observation $\mathbf{O}_{t}$ as input, predicts $T_{p}$ action sequence $\mathbf{A}_{t}$ and executes $T_{a}$ action sequence. The action sequence generation process is a conditional denoising diffusion process modeled by $p_{\theta}(\mathbf{A}_{t}|\mathbf{O}_{t})$, where $\mathbf{A}_{t}\in\mathbb{R}^{T_{p}\times D_{a}}$ and $\mathbf{O}_{t}\in\mathbb{R}^{T_{o}\times D_{o}}$, $D_{a}$ and $D_{o}$ represent the action and observation’s dimension respectively. Specifically, starting from a pure Gaussian noise sample, diffusion policy leverage the noise prediction network $\varepsilon_{\theta}$ to predict and remove noise at each denoising step, iterating for $K$ steps to generate a clean sample $\mathbf{A}_{t}^{0}$. The action sequence generation process is described as equation 3, where $\mathbf{A}_{t}^{k}$ is a partial denoising action in the timestep $t$ with noise level $k$, $\mathbf{Z}$ is a standard Gaussian random variable.
@@
The training loss is:
$L(\theta):=\mathbb{E}_{k,\mathbf{A}_{t}^{0},\boldsymbol{\epsilon}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\mathbf{O}_{t},\sqrt{\bar{\alpha}_{k}}\mathbf{A}_{t}^{0}+\sqrt{1-\bar{\alpha}_{k}}\boldsymbol{\epsilon},k\right)\right\|^{2}\right]$ (4)
We write $[a,b]$ to denote the set $\{a,a+1,\cdots,b\}$ and $[a,b)$ to denote the set $\{a,a+1,\cdots,b-1\}$. We write $\boldsymbol{x}_{a:b}$ to denote the set $\{\boldsymbol{x}_{i}:I\in[a,b)\}$ and $[K]$ to denote the set $\{1,\cdots,K\}$. We define $\mathbf{A}_{t}=[\boldsymbol{a}_{t:t+T_{a}},\boldsymbol{\tilde{a}}_{t+T_{a}:t+T_{p}}]$, where $\boldsymbol{a}_{t:t+T_{a}}$ is the executed part and $\boldsymbol{\tilde{a}}_{t+T_{a}:t+T_{p}}$ is the unexecuted part in timestep $t$.
@@ 3.2 Thresholding Mechanism
To determine which partial denoised action from historical observations should be used as the initialization for denoising at the current timestep, we employ Tweedie’s approach (Efron, 2011; Chung et al., 2022; Kim & Ye, 2021). Given that the forward process of diffusion policy follows
@@
we derive the posterior expectation $\mathbb{E}[\boldsymbol{a}_{\tau:\tau-T_{a}+T_{p}}^{0}\mid\mathbf{O}_{t},\boldsymbol{a}_{\tau:\tau-T_{a}+T_{p}}^{k}]$ which serves as the one-step estimation of the partial denoised actions $\boldsymbol{a}_{\tau:\tau-T_{a}+T_{p}}^{k_{i}}$, denoted by $\hat{\boldsymbol{a}}_{\tau:\tau-T_{a}+T_{p}}^{k_{i}}$. This estimation is conditioned on the current observation $\mathbf{O}_{t}$, as formulated in Proposition 6. Here $\tau<t$ represents a historical decision step and $k$ denotes the noise level of the partial denoised action $\boldsymbol{a}_{\tau:\tau-T_{a}+T_{p}}^{k}$.
@@
The dependency between past and current actions is measured by the Euclidean distance $\|\hat{\boldsymbol{a}}_{\tau:\tau-T_{a}+T_{p}}^{k_{i}}-\boldsymbol{\tilde{a}}_{t:t-T_{a}+T_{p}}\|_{2}$ between $\hat{\boldsymbol{a}}_{\tau:\tau-T_{a}+T_{p}}^{k_{i}}$ and the reference action $\boldsymbol{\tilde{a}}_{t:t-T_{a}+T_{p}}$. If this distance falls below a predefined threshold $\epsilon$, the partial denoised action $\boldsymbol{a}_{\tau:\tau+T_{p}}^{k}$ is selected, as it is likely to converge to the desired action $\boldsymbol{a}_{t:t+T_{p}}$.
@@ 3.3 Implementation Details
Falcon can also be compatible with other diffusion policy acceleration techniques by replacing the solver in Line 7 and Line 18 with other SDE/ODE solvers like DDIM (Song et al., 2020a) and DPMSolver (Lu et al., 2022). The detailed pseudocode is in the Appendix A.
@@ 4 Experiments
Our experiments aim to address three key questions: (1) Can Falcon accelerate diffusion policy, and does it further enhance speed when integrated with other acceleration algorithms (Section 4.3)? (2) Does Falcon maintain its acceleration advantage in long-sequence tasks (Section 4.4)? (3) Can Falcon retain the ability to express multimodality while achieving speed improvements (Section 4.5)?
4.1 Metrics
4.2 Simulation Environments and datasets
Robomimic (Mandlekar et al., 2021) is a large-scale benchmark designed to evaluate robotic manipulation algorithms using human demonstration datasets. Each task provides two types of human demonstrations: proficient human (PH) demonstrations, representing expert performance, and mixed proficient and nonproficient human (MH) demonstrations. Each environment uses an action prediction horizon $T_{p}=16$ and an action execution horizon $T_{a}=8$, and all tasks use state-based observations. We construct a CNN-based Diffusion Policy (Chi et al., 2023) with DDPM scheduler using 100 denoising steps and DDIM/DPMSolver scheduler using 16 denoising steps.
Franka Kitchen (Gupta et al., 2019) is designed for evaluating algorithms on long-horizon, multi-stage robotics tasks. The environment features a 9-DOF Franka robot and a kitchen scene with 7 objects for interaction and comes with a human demonstration dataset of 566 demonstrations, each completing 4 tasks in arbitrary order. The task involves action sequences of dimension 112 and an episode length of 1200, with an action prediction horizon of $T_{p}=16$ and an action execution horizon of $T_{a}=8$. We construct a transformer-based Diffusion Policy using a DDPM scheduler with 100-step discretization.
MetaWorld (Yu et al., 2020) is an open-source simulation benchmark for meta-reinforcement learning and multi-task learning. It consists of 50 different robotic manipulation tasks categorized into different difficulty levels ranging from simple to very challenging.
Lift Can Square Transport ToolHang
ph mh ph mh ph mh ph mh ph
DDIM+Falcon $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $1.00_{\pm 1.00}$ $0.98_{\pm 0.14}$ $0.91_{\pm 0.28}$ $0.85_{\pm 0.36}$ $0.74_{\pm 0.44}$ $0.63_{\pm 0.48}$ $0.51_{\pm 0.49}$
Table 1: Success Rate in Robomimic. We present the success rate with 200 evaluation episodes in the format of (mean of success rate) $\pm$ (standard deviation of success rate). Our results show that Falcon matches the original methods
@@ 4.2 Simulation Environments and datasets
DDIM+Falcon $\mathbf{7.5_{\pm 0.2}}$ $\mathbf{7.7_{\pm 0.2}}$ $\mathbf{6.5_{\pm 0.3}}$ $\mathbf{7.6_{\pm 2.1}}$ $\mathbf{7.6_{\pm 1.0}}$ $\mathbf{7.2_{\pm 1.2}}$ $\mathbf{9.2_{\pm 0.6}}$ $\mathbf{9.0_{\pm 1.1}}$ $\mathbf{8.4_{\pm 0.5}}$
@@
Table 2: NFE in Robomimic. We present the NFE with 200 evaluation episodes in the format of (mean of NFE) $\pm$ (standard deviation of NFE). Our results show that Falcon drastically reduces the denoising steps compared with the original methods

### v2

# Haojun Chen    Minghao Liu    Chengdong Ma    Xiaojian Ma    Zailin Ma    Huimin Wu    Yuanpei Chen    Yifan Zhong    Mingzhi Wang    Qing Li    Yaodong Yang

@@ Abstract
Diffusion policies are widely adopted in complex visuomotor tasks for their ability to capture multimodal action distributions. However, the multiple sampling steps required for action generation significantly harm real-time inference efficiency, which limits their applicability in real-time decision-making scenarios. Existing acceleration techniques either require retraining or degrade performance under low sampling steps. Here we propose Falcon, which mitigates this speed-performance trade-off and achieves further acceleration. The core insight is that visuomotor tasks exhibit sequential dependencies between actions. Falcon leverages this by reusing partially denoised actions from historical information rather than sampling from Gaussian noise at each step. By integrating current observations, Falcon reduces sampling steps while preserving performance. Importantly, Falcon is a training-free algorithm that can be applied as a plug-in to further improve decision efficiency on top of existing acceleration techniques. We validated Falcon in 48 simulated environments and 2 real-world robot experiments. demonstrating a 2-7x speedup with negligible performance degradation, offering a promising direction for efficient visuomotor policy design.
@@ 1 Introduction
Diffusion policies have demonstrated remarkable success in addressing complex visuomotor tasks in robotics (Chi et al., 2023; Reuss et al., 2023; Ze et al., 2024b, a; Ravan et al., 2024; Yang et al., ), thanks to their ability to model complex multimodal distributions and maintain stable training dynamics. Essentially, diffusion policies rely on the reverse sampling process of a stochastic differential equation (SDE) (Song et al., b; Karras et al., 2022), where actions are iteratively sampled starting from a standard normal distribution. Each sampling step involves drawing a sample from a Brownian motion distribution, incrementally denoising the initial sample to generate the final action. However, the iterative sampling process required for action generation makes these methods computationally slow, particularly when applied to sequential decision-making in real-time environments (Chen et al., b; Janner et al., 2022; Wang et al., b; Yang et al., 2023; Hansen-Estruch et al., 2023; Chen et al., a).
To address these challenges, existing work (Song et al., a, b; Lu et al., 2022; Zhang & Chen, ) reformulates the sampling process as an ordinary differential equation (ODE) and uses numerical solvers to reduce the number of denoising steps (Song et al., a; Lu et al., 2022; Zhang & Chen, ). However, these solvers often suffer from approximation errors when using very few steps, degrading policy performance (Zhao et al., 2024; Wang et al., a). Alternatively, distillation-based approaches (Salimans & Ho, ; Song et al., 2023; Kim et al., ; Prasad et al., 2024; Wang et al., 2024) accelerate inference via one-step generation. However, they typically require task-specific retraining and may degrade multimodal expressiveness (Prasad et al., 2024), making them less flexible and unsuitable as plug-and-play solutions for diverse tasks. Another related approach is Streaming Diffusion Policy (SDP) (Høeg et al., 2024), which also uses partial denoising for acceleration but requires task-specific retraining and a large noise-time buffer, limiting flexibility and memory efficiency.
@@
In this work, we introduce Falcon (Fast visuomotor policies via partial denoising), a novel approach designed to bridge the gap between acceleration and performance preservation in diffusion-based visuomotor policies. The key insight behind Falcon is that sequential dependencies in visuomotor tasks can be exploited to accelerate action generation while maintaining multimodal expressiveness. To effectively leverage this property, we first utilize the previous action prediction as the reference action, then introduce a thresholding mechanism combined with one-step estimation to evaluate which partial denoising actions are dependent on current timesteps in parallel. Using a temperature-scaled softmax, we then select the most suitable partial denoising action to continue the sampling process, preserving both performance and efficiency. By avoiding the conventional practice of starting the denoising process from a standard normal distribution at each decision step, Falcon begins from partial denoised actions derived from historical observations, significantly reducing the number of sampling steps required. Importantly, Falcon is a training-free method, which allows it to be applied as a plug-in module, enhancing the efficiency of existing diffusion-based policies without additional training or extensive modifications. It integrates seamlessly with solvers like DDIM (Song et al., a) and DPMSolver (Lu et al., 2022) to further enhance acceleration.
We evaluate Falcon in both simulated and real-world environments. In simulation, we test Falcon across 48 tasks spanning five widely used benchmarks, including RoboMimic (Mandlekar et al., 2022), RoboSuite Kitchen (Gupta et al., 2020), BlockPush (Shafiullah et al., 2022), MetaWorld (Yu et al., 2020) and Maniskill2 (Gu et al., 2023). For real-world experiments, we deploy Falcon on a dual-arm robot platform in two manipulation tasks—dexterous grasping and high-precision insertion. In both settings, Falcon achieves a 2–7× acceleration in action sampling while preserving high task success rates, validating its effectiveness and practical applicability.
@@
In summary, this work makes three main contributions: First, we introduce Falcon, which leverages sequential dependencies in visuomotor tasks to perform partial denoising, significantly reducing the number of sampling steps while preserving the ability to model multimodal action distributions. Second, Falcon functions as a training-free plug-in that enhances existing diffusion-based policies, integrating seamlessly with solvers like DDIM and DPMSolver to further accelerate action generation. Third, we demonstrate Falcon’s effectiveness across 48 simulated environments and 2 real-world robot tasks, achieving a 2-7× speedup in inference while maintaining high action quality.
@@ 2 Preliminaries
In this section, we briefly introduce the diffusion models, how the diffusion models are used for diffusion policies in visuomotor tasks, the acceleration techniques in the diffusion-based models, and terminology.
Diffusion Models.   Diffusion models, such as DDPM (Ho et al., 2020), are generative models that learn data distributions by progressively corrupting data through noise in a forward process and then reconstructing it in a reverse denoising process. In the forward process, a data point $\boldsymbol{x}_{0}$ is corrupted K timesteps, resulting in:
@@
where $\mu_{\theta}$ is the predicted mean, $\sigma_{t}^{2}$ is fixed according to the forward process, $\alpha_{i}$ is the variance schedule in the forward diffusion process.
Diffusion Policies.   Diffusion policy (Chi et al., 2023) extends diffusion models as a powerful policy for visuomotor tasks. At timestep $t$, diffusion policy takes the latest $T_{o}$ steps of observation $\mathbf{O}_{t}$ as input, predicts $T_{p}$ action sequence $\mathbf{A}_{t}$ and executes $T_{a}$ action sequence. The action sequence generation process is a conditional denoising diffusion process modeled by $p_{\theta}(\mathbf{A}_{t}|\mathbf{O}_{t})$, where $\mathbf{A}_{t}\in\mathbb{R}^{T_{p}\times D_{a}}$ and $\mathbf{O}_{t}\in\mathbb{R}^{T_{o}\times D_{o}}$, $D_{a}$ and $D_{o}$ represent the action and observation’s dimension respectively. Specifically, starting from a pure Gaussian noise sample, diffusion policy leverage the noise prediction network $\varepsilon_{\theta}$ to predict and remove noise at each denoising step, iterating for $K$ steps to generate a clean sample $\mathbf{A}_{t}^{0}$. The action sequence generation process is described as Eq. 3, where $\mathbf{A}_{t}^{k}$ is a partial denoising action in the timestep $t$ with noise level $k$, $\mathbf{Z}$ is a standard Gaussian random variable.
@@
The training loss is
$L(\theta)$ $=\mathbb{E}_{k,\mathbf{A}_{t}^{0},\boldsymbol{\epsilon}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\mathbf{O}_{t},\mathbf{A}_{t}^{k},k\right)\right\|^{2}\right]$ (4)
$\mathbf{A}_{t}^{k}$ $=\sqrt{\bar{\alpha}_{k}}\,\mathbf{A}_{t}^{0}+\sqrt{1-\bar{\alpha}_{k}}\,\boldsymbol{\epsilon}$
Acceleration techniques in the diffusion-based models.   Some popular works (Song et al., b; Karras et al., 2022; Song et al., a; Lu et al., 2022) interpret the diffusion model as an ODE. Specifically, DDIM (Song et al., a) generalizes the Markov forward process in DDPM to a non-Markov process, allowing the use of shorter Markov chains during sampling. The iterative steps of its sampling process can be rewritten in a form similar to Euler integration, which is a discrete solution process for a specific ODE. DPMSolver (Lu et al., 2022) accelerates sampling by analytically computing the linear part of the diffusion ODE solution and using an exponential integrator to approximate the nonlinear part. But these methods can reduce the quality degradation of few-step sampling (Shih et al., 2024). SDP (Høeg et al., 2024) improves the sampling speed by generating partial denoised actions with different noise levels, but relies on handcrafted noise schedules and large buffers, limiting its flexibility. Consistency Policy (CP) (Prasad et al., 2024) distills a pretrained diffusion policy into a one-step sampler via self-consistency on its probability flow ODE. While fast, this requires retraining and may compromise multimodality.
We write $[a,b)$ to denote the set $\{a,a+1,\cdots,b-1\}$, $\boldsymbol{x}_{a:b}$ to denote the set $\{\boldsymbol{x}_{i}:I\in[a,b)\}$ and $[K]$ to denote the set $\{1,\cdots,K\}$. We define $\mathbf{A}_{t}=[\boldsymbol{a}_{t:t+T_{a}},\boldsymbol{\tilde{a}}_{t+T_{a}:t+T_{p}}]$, where $\boldsymbol{a}_{t:t+T_{a}}$ is the executed part and $\boldsymbol{\tilde{a}}_{t+T_{a}:t+T_{p}}$ is the unexecuted part in timestep $t$.
@@ 3.2 Thresholding Mechanism
To determine which partial denoised action from historical observations should be used as the initialization for denoising at the current timestep, we employ Tweedie’s approach (Efron, 2011; Chung et al., ; Kim & Ye, 2021). Given that the forward process of diffusion policy follows
@@
We derive the posterior expectation, which serves as the one-step estimation of the partial denoised actions $\boldsymbol{a}_{\tau:\tau-T_{a}+T_{p}}^{k_{i}}$, denoted by $\hat{\boldsymbol{a}}_{\tau:\tau-T_{a}+T_{p}}^{k_{i}}$:
$\hat{\boldsymbol{a}}_{\tau:\tau-T_{a}+T_{p}}^{k_{i}}=\mathbb{E}[\boldsymbol{a}_{\tau:\tau-T_{a}+T_{p}}^{0}\mid\mathbf{O}_{t},\boldsymbol{a}_{\tau:\tau-T_{a}+T_{p}}^{k}].$ (6)
This estimation is conditioned on the current observation $\mathbf{O}_{t}$, as formulated in Proposition 7. Here, $\tau<t$ represents a historical decision step and $k$ denotes the noise level of the partial denoised action.
@@
The dependency between past and current actions is measured by the Euclidean $\|\hat{\boldsymbol{a}}_{\tau:\tau-T_{a}+T_{p}}^{k_{i}}-\boldsymbol{\tilde{a}}_{t:t-T_{a}+T_{p}}\|_{2}$ between $\hat{\boldsymbol{a}}_{\tau:\tau-T_{a}+T_{p}}^{k_{i}}$ and the reference action $\boldsymbol{\tilde{a}}_{t:t-T_{a}+T_{p}}$. If this distance falls below a predefined threshold $\epsilon$, the partial denoised action $\boldsymbol{a}_{\tau:\tau+T_{p}}^{k}$ is selected, as it is likely to converge to the desired action $\boldsymbol{a}_{t:t+T_{p}}$.
@@
Lift Can Square Transport ToolHang
ph mh ph mh ph mh ph mh ph
SDP(DDPM) $1.00_{\pm 0.00}$ $0.98_{\pm 0.12}$ $0.97_{\pm 0.17}$ $0.94_{\pm 0.24}$ $0.88_{\pm 0.32}$ $0.80_{\pm 0.40}$ $0.81_{\pm 0.39}$ $0.46_{\pm 0.49}$ $0.58_{\pm 0.50}$
DDIM+Falcon $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $1.00_{\pm 1.00}$ $0.98_{\pm 0.14}$ $0.91_{\pm 0.28}$ $0.85_{\pm 0.36}$ $0.81_{\pm 0.39}$ $0.63_{\pm 0.48}$ $0.54_{\pm 0.50}$
Table 1: Success Rate in Robomimic. We present the success rate with 200 evaluation episodes in the format of (mean of success rate) $\pm$ (standard deviation of success rate). Our results show that Falcon matches the original methods.
@@ 3.3 Implementation Details
Falcon can also be compatible with other diffusion policy acceleration techniques by replacing the solver in Line 7 and Line 18 with other SDE/ODE solvers like DDIM (Song et al., a) and DPMSolver (Lu et al., 2022). The detailed pseudocode is in the Appendix B.
@@
Figure 3: Multimodal behavior. At the given symmetric configuration in the PushT task, DDPM+Falcon (right) preserves multimodal behaviors by producing diverse reaching trajectories. In contrast, the Consistency Policy (left) shows bias toward one mode. Each line represents one rollout trajectory, plotted from 20 trajectories sampled at the same state.
@@ 4 Experiments
Our experiments aim to address four key questions: (1) Can Falcon accelerate diffusion policy, and does it further enhance speed when integrated with other acceleration algorithms (Section 4.3)? (2) Can Falcon effectively accelerate diffusion policies in real-world robotic settings while preserving performance (Section 4.5)? (3) Does Falcon maintain its acceleration advantage in long-sequence tasks (Section E.3)? (4) Can Falcon retain the ability to express multimodality while achieving speed improvements (Section 4.6)?
@@
DDIM+Falcon $\mathbf{7.5_{\pm 0.2}}$ $\mathbf{7.7_{\pm 0.2}}$ $\mathbf{6.5_{\pm 0.3}}$ $\mathbf{7.6_{\pm 2.1}}$ $\mathbf{7.6_{\pm 1.0}}$ $\mathbf{7.2_{\pm 1.2}}$ $\mathbf{10.5_{\pm 2.0}}$ $\mathbf{9.0_{\pm 1.1}}$ $\mathbf{10.1_{\pm 1.4}}$
@@
Table 2: NFE in Robomimic. We present the NFE with 200 evaluation episodes in the format of (mean of NFE) $\pm$ (standard deviation of NFE). Our results show that Falcon drastically reduces the denoising steps compared with the original methods.
4.1 Metrics
Lift Can Square Transport ToolHang
ph mh ph mh ph mh ph mh ph
DPMSolver* $1.00_{\pm 0.00}$ $0.95_{\pm 0.20}$ $0.94_{\pm 0.23}$ $0.96_{\pm 0.20}$ $0.84_{\pm 0.37}$ ($\downarrow$10.6%) $0.58_{\pm 0.50}$ ($\downarrow$35.6%) $0.69_{\pm 0.46}$ ($\downarrow$6.8%) $0.54_{\pm 0.50}$ $0.53_{\pm 0.50}$
[…]
Figure 4: Left: Ablation analysis on threshold $\epsilon$. The effect of varying the $\epsilon$ threshold on both the score (red circles) and speedup (blue squares), demonstrating the trade-off between performance and speed as $\epsilon$ changes. Right: Ablation analysis on exploration rate $\delta$. The score (red circles) increases sharply with $\delta$ and then stabilizes, while the speedup (blue squares) decreases as the exploration rate rises. This study highlights the balance between exploration and exploitation.
4.2 Simulation and Real-World Environments
We evaluate Falcon across diverse simulated and real-world environments to demonstrate its generality and effectiveness. Simulation benchmarks include RoboMimic (Mandlekar et al., 2022), Franka Kitchen (Gupta et al., 2020), BlockPush (Shafiullah et al., 2022), PushT (Chi et al., 2023), MetaWorld (Yu et al., 2020), and ManiSkill2 (Gu et al., 2023), covering short- and long-horizon tasks, as well as multimodal behavior tasks. For real-world evaluation, we deploy Falcon on a dual-arm robot platform in two manipulation tasks: dexterous grasping and high-precision insertion. Details of simulation settings, network architectures, and hardware platforms are provided in the Appendix A.